# Cartesian subgroups in graph products of groups

### v1

@@ Abstract.
The kernel of the natural projection of a graph product of groups onto their direct product is called the Cartesian subgroup. This construction generalises commutator subgroups of right-angled Coxeter and Artin groups. Using theory of polyhedral products, we give a lower and an upper bound on the number of relations in presentations of Cartesian groups and on their deficiency. The bounds are related to the fundamental groups of full subcomplexes in the clique complex, and the lower bound coincide with the upper bound if these fundamental groups are free or free abelian.
@@ 1. Introduction
$\operatorname{Cart}(\underline{G},\Gamma):=\operatorname{Ker}(\underline{G}^{\Gamma}\to G_{i}\times\dots\times G_{m})$
@@
The generators are explicit (each generator is of the form $L_{\underline{g}}(i,J):=\prod_{j\in J}g_{j}\cdot g_{i}^{-1}\cdot(\prod_{j\in J\setminus i}g_{j})^{-1}$), while the relations can be computed by a recursive algorithm (see Theorem 4.11). Moreover, the generators and relations have a geometric meaning in terms of paths in certain polyhedral products, and this result can be extended to the case of infinite graph products (see Subsection 6.1). We also give a homological lower bound on the size of a presentation:
@@
As a special case we obtain a small presentation of the commutator subgroup $\mathrm{RC}_{\mathcal{K}}^{\prime}$ in any right-angled Coxeter group, and a lower bound on the number of relations in any its presentation:
@@
Acknowledgements and funding
This work was carried out within the project “Mirror Laboratories” of HSE University, Russian Federation, and is supported by the Russian Science Foundation under grant no. 24-71-00059, http://rscf.ru/en/project/24-71-00059/.
The author woud like to thank Taras E. Panov for suggesting the problem and valuable advice, and Li Cai, Anton A. Klyachko, Temurbek Rahmatullaev and Andrei Yu. Vesnin for helpful discussions.
@@ 2.1. Simplicial complexes
Every face $I\in\mathcal{K}$ corresponds to a geometric face $|I|=\mathrm{conv}\{e_{i}:~{}i\in I\}$ of the standard geometric simplex $L^{m-1}=\mathrm{conv}\{e_{i}:i\in[m]\}\subset\mathbb{R}^{m}.$ The union of these geometric faces is called the geometric realization $|\mathcal{K}|$ of a complex $\mathcal{K}.$
@@ 2.2. Polyhedral products
$H_{i}(\mathcal{R}_{\mathcal{K}})\cong\bigoplus_{J\subset[m]}{\widetilde{H}}_{i-1}(\mathcal{K}_{J}).$
A far-reaching generalization of this result is the following stable homotopy decomposition for polyhedral products of the form $(\operatorname{cone}\underline{A},\underline{A})^{\mathcal{K}}:$
@@
Under the conditions of Theorem 2.2, we have
@@ 2.3. Classifying spaces of graph products and Cartesian subgroups
Davis described the fundamental groups $\pi_{1}((\underline{X},\underline{A})^{\mathcal{K}})$ of general polyhedral products algebraically as relative graph products [Dav12, Theorem 2.18].
@@
An almost complete answer to the question “Which polyhedral products are aspherical?” was obtained by Davis and Kropholler [DK15, Theorem 1].
@@ 4.1. Generators
For each $J\subset[m]$, choose a subset $\Theta(J)\subset J\setminus\{\max(J)\}$ such that $\Theta(J)\sqcup\{\max(J)\}$ contains exactly one vertex from each path component of $\mathcal{K}_{J}$.
@@
Panov–Veryovkin generators are important for the analogy between the group $\pi_{*}(\mathcal{R}_{\mathcal{K}})$ and the Hopf algebra $H_{*}(\Omega\mathcal{Z}_{\mathcal{K}};\mathbb{Z})$ (see Subsection 6.4 below). However, our choice of generators is more suitable for computations.
@@
For right-angled Coxeter groups $\underline{G}^{\mathcal{K}}=(\underline{\mathbb{Z}_{2}})^{\mathcal{K}}=\mathrm{RC}_{\mathcal{K}}$ we omit the subscript ${\underline{g}}$ since then $|G_{j}^{*}|=1$. Li Cai proved ([Cai21], unpublished) that the set $\widehat{D}=\{\widehat{L}(i,J):J\subset[m],i\in\Theta(J)\}\subset\mathrm{RC}_{\mathcal{K}}^{\prime}$ generates the group $\mathrm{RC}_{\mathcal{K}}^{\prime}.$ His approach relies on a study of the Davis complex [Dav08, Chapter 7], a contractible cubical complex which is the universal covering of $\mathcal{R}_{\mathcal{K}}.$
@@ 4.2. Reduction to the distinguished elements
    in $\mathcal{K}_{J},$ i.e. a sequence of edges $\{i_{t},i_{t+1}\}\in\mathcal{K}_{J}$. Applying the relation (4.1), we replace $L_{\underline{g}}(i_{t},J)$ with a product of $L_{\underline{g}}(i_{t+1},J)$ and generators of smaller length. On the last step we replace $L_{\underline{g}}(\max(J),J)$ with $1$ (using (4.2)). Therefore, $L_{\underline{g}}(i,J)$ is replaced with a word on elements of smaller length. By the inductive hypothesis, each of this elements can be expressed through the distinguished elements.
@@ 4.4. Examples
$\cdot$ $L(4,1235)L(3,1235)^{-1}\cdot L(5,1235)L(4,1234)^{-1}\cdot L(1,1234)L(5,2345)^{-1}\Big{)}=$
@@ 5.1. Outline of the proof
Consider the polyhedral product $B:=(\operatorname{cone}(\underline{G}),\underline{G})^{\mathcal{K}}\simeq(E\underline{G},\underline{G})^{\mathcal{K}}$. In Proposition 5.11 we give a geometric interpretation of the isomorphism $\pi_{1}(B)\cong\operatorname{Cart}(\underline{G},\mathcal{K})$, identifying certain loops in $B$ with elements of $\operatorname{Cart}(\underline{G},\mathcal{K})$. Then, following Li Cai [Cai21], we represent $B$ as the union of contractible subspaces ${\mathrm{cc}}(\mathcal{K})_{\underline{g}}$ (see (5.1)), ${\underline{g}}\in\prod_{i=1}^{m}G_{i}$.
For a subset $Q\subset\prod_{i=1}^{m}G_{i}$, denote $B_{Q}:=\bigcup_{{\underline{h}}\in Q}{\mathrm{cc}}(\mathcal{K})_{\underline{h}}$. Then $B_{Q\sqcup\{{\underline{g}}\}}=B_{Q}\cup{\mathrm{cc}}(\mathcal{K})_{\underline{g}},$ where ${\mathrm{cc}}(\mathcal{K})_{\underline{g}}$ is contractible. In some cases the intersection $B_{Q}\cap{\mathrm{cc}}(\mathcal{K})_{\underline{g}}$ is homotopy equivalent to $|\mathcal{K}_{J}|$ for $J=\operatorname{supp}{\underline{g}}\subset[m]$ (Proposition 5.9). Hence, by an application of the van Kampen theorem (Lemma 5.1), the group $\pi_{1}(B_{Q\sqcup\{{\underline{g}}\}})$ is obtained from $\pi_{1}(B_{Q})$ by imposing $\operatorname{rank}\Pi_{1}(\mathcal{K}_{J})$ relations and then adding $\widetilde{b}_{0}(\mathcal{K}_{J})$ generators. By induction, in Theorem 5.14 we give presentations for fundamental groups of $\pi_{1}(B_{Q})$ whenever $Q\subset\prod_{i=1}^{m}G_{i}$ is an admissible subset (see Definition 5.6). A presentation of $\pi_{1}(B)$ is then obtained by passing to the limit (Lemma 5.21).
@@ 5.2. The geometric model for the classifying space
The polyhedral product $(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$ and its subspace ${\mathrm{cc}}(\mathcal{K})_{\underline{1}}$ is shown in Figure 2 for the case $\mathcal{K}=\bullet\bullet\!\!-\bullet$ and $\underline{G}=(\mathbb{Z}_{2},\mathbb{Z}_{2},\mathbb{Z}_{2})$ (left) and $\underline{G}=(\mathbb{Z}_{2},\mathbb{Z}_{3},\mathbb{Z}_{2})$ (right).
@@
Let $x\in X$ be a basepoint. For each $\alpha\in PC,$ choose a basepoint $a_{\alpha}\in A_{\alpha}$, a path $t_{\alpha}$ in $X$ from $x$ to $a_{\alpha}$, and a set of loops $\{\lambda_{i,\alpha}:i\in I_{\alpha}\}$ in $A_{\alpha}$ that generate the group $\pi_{1}(A_{\alpha},a_{\alpha}).$
@@
Now denote $T:=\overline{t_{0}}\cup\bigcup_{\alpha\in PC\setminus\{0\}}\overline{s_{\alpha}}.$ We have the following pushout diagram of inclusion of connected CW-complexes with the basepoint $x$:
${A\cup T\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i}$${X\cup T\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C\cup T\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Y.}$
@@
    The isomorphism $\Pi_{1}(A)\overset{\cong}{\longrightarrow}\pi_{1}(A\cup T,x)$ is described as follows: if $\lambda$ is a loop in $A_{\alpha}$ with the basepoint $a_{\alpha},$ then $t_{\alpha}\cdot\lambda\cdot t_{\alpha}^{-1}$ is the corresponding loop in $A\cup T$ with the basepoint $x.$
  • The space $C\cup T$ is obtained from a simply connected space $C$ by attaching a segment $[a_{0},x]$ at the point $a_{0}\in C$, hence $\pi_{1}(C\cup T,x)$ is trivial.
  • The space $X\cup T$ is obtained from $X$ by attaching 1-cells $\{s_{\alpha}:\alpha\in PC\setminus\{0\}\}$ connecting the points $a_{0}$ and $a_{\alpha}.$ Clearly, $X\cup T\simeq(X\cup T)/T\simeq X\vee\bigvee_{\alpha\in PC\setminus\{0\}}S^{1}.$ Hence $\pi_{1}(X\cup T)$ is the free product of $\pi_{1}(X)$ with the free group on $\widetilde{b}_{0}(A)$ generators,
    $\pi_{1}(X\cup T,x)\cong\pi_{1}(X,x)\ast F(\{t_{0}\cdot s_{\alpha}\cdot t_{\alpha}^{-1}:\alpha\in PC\setminus\{0\}\}).$
@@
where $i_{*}$ is the free product of maps $\pi_{1}(A,a_{\alpha})\to\pi_{1}(X),$ $\lambda\mapsto t_{\alpha}\cdot\lambda\cdot t_{\alpha}^{-1}.$
@@ 5.3. Subspaces of ${\mathrm{cc}}(\mathcal{K})$
By the previous lemma, we have ${\mathrm{cc}}(\mathcal{K})\cap\{x_{j}=0\}=\iota_{\mathcal{K}}(|\operatorname{st}_{\mathcal{K}^{\prime}}(\{j\})|)$, so we obtain the first statement by taking the union over $j\in J$.
@@
Consider the full subcomplex of $\mathcal{K}^{\prime}$ on the vertex set $\{(I):I\in\mathcal{K},I\cap J\neq\varnothing\}$. Since $I\cap J\neq\varnothing$ if and only if $\{j\}\subsetneq I$ for some $j\in J$, this full subcomplex coincides with $\mathcal{N}$.
@@
Hence we can define a strict deformation retraction of $|\mathcal{N}|$ onto $\beta(|\mathcal{K}_{J}|)$ as follows: for every $I\in\mathcal{K}$ such that $I\cap J\neq\varnothing,$ we move $b_{I}$ linearly to the point $b_{I\cap J}$ along the segment $|(I\cap J\subsetneq I)|\subset|\mathcal{N}|;$ we extend this homotopy linearly onto geometric simplices of $|\mathcal{N}|$. This homotopy is well defined: if $(I_{1}\subsetneq\dots\subsetneq I_{r})\in\mathcal{N}$, then the geometric simplex $\operatorname{conv}(b_{I_{1}},\dots,b_{I_{r}})\subset|\mathcal{N}|$ is linearly deformed onto the geometric simplex $\operatorname{conv}(b_{I_{1}\cap J},\dots,b_{I_{r}\cap J})$; the image of the homotopy lies in $\operatorname{conv}(b_{I_{1}},\dots,b_{I_{r}},b_{I_{1}\cap J},\dots,b_{I_{r}\cap J})\subset\beta(|I_{r}|)\subset|\mathcal{N}|.$

### v2

@@ Abstract.
The kernel of the natural projection of a graph product of groups onto their direct product is called the Cartesian subgroup of the graph product. This construction generalises commutator subgroups of right-angled Coxeter and Artin groups. Using theory of polyhedral products, we give a lower and an upper bound on the number of relations in presentations of Cartesian groups and on their deficiency. The bounds are related to the fundamental groups of full subcomplexes in the clique complex, and the lower bound coincide with the upper bound if these fundamental groups are free or free abelian.
@@ 1. Introduction
$\operatorname{Cart}(\underline{G},\Gamma):=\operatorname{Ker}(\underline{G}^{\Gamma}\to G_{1}\times\dots\times G_{m})$
@@
We provide an alternative minimal set of generators for $\operatorname{Cart}(\underline{G},\mathcal{K}),$ following the approach of Li Cai [Cai21*], and describe a small sufficient set of relations between them. Let $X=\bigsqcup_{\alpha}X_{\alpha}$ be a topological space with path components $\{X_{\alpha}\}$. Consider the group
@@
See Theorem 4.11 for a detailed description of this presentation. The generators are explicit (each generator is of the form $L_{\underline{g}}(i,J):=\prod_{j\in J}g_{j}\cdot g_{i}^{-1}\cdot(\prod_{j\in J\setminus i}g_{j})^{-1}$), while the relations can be computed by a recursive algorithm. Moreover, the generators and relations have a geometric meaning in terms of paths in certain polyhedral products, and this result can be extended to the case of infinite graph products (see Subsection 6.1). We also give a homological lower bound on the size of a presentation:
@@
As a special case we obtain a small presentation of the commutator subgroup $\mathrm{RC}_{\mathcal{K}}^{\prime}$ in any right-angled Coxeter group, and a lower bound on the number of relations in any of its presentations:
@@
Acknowledgements
The author woud like to thank Taras E. Panov for suggesting the problem and valuable advice, Li Cai, Anton A. Klyachko, Temurbek Rahmatullaev and Andrei Yu. Vesnin for helpful discussions, and the anonymous referee for careful reading of the text and important corrections.
@@ 2.1. Simplicial complexes
Every face $I\in\mathcal{K}$ corresponds to a geometric face $|I|=\mathrm{conv}\{e_{i}:~{}i\in I\}$ of the standard geometric simplex $\Delta^{m-1}=\mathrm{conv}\{e_{i}:i\in[m]\}\subset\mathbb{R}^{m}.$ The union of these geometric faces is called the geometric realization $|\mathcal{K}|$ of a complex $\mathcal{K}.$
@@ 2.2. Polyhedral products
For any $i\geq 0$ and any group of coefficients, we have $H_{i}(\mathcal{R}_{\mathcal{K}})\cong\bigoplus_{J\subset[m]}{\widetilde{H}}_{i-1}(\mathcal{K}_{J}).$
(For $i=0$ we obtain $H_{0}(\mathcal{R}_{\mathcal{K}};\mathbf{k})\cong{\widetilde{H}}_{-1}(\varnothing;\mathbf{k})\cong\mathbf{k}$.) A far-reaching generalization of this result is the following stable homotopy decomposition for polyhedral products of the form $(\operatorname{cone}\underline{A},\underline{A})^{\mathcal{K}}:$
@@
Under the conditions of Theorem 2.2, we have, for any $i\geq 0$,
@@ 2.3. Classifying spaces of graph products and Cartesian subgroups
Davis described the fundamental groups $\pi_{1}((\underline{X},\underline{A})^{\mathcal{K}})$ of general polyhedral products algebraically as relative graph products [Dav12, Theorem 2.18]. An almost complete answer to the question “Which polyhedral products are aspherical?” was obtained by Davis and Kropholler [DK15, Theorem 1].
@@ 4.1. Generators
For each $J\subset[m]$, choose a subset $\Theta(J)\subset J\setminus\{\max(J)\}$ such that $\Theta(J)\sqcup\{\max(J)\}$ contains exactly one vertex from each path component of $|\mathcal{K}_{J}|$.
@@
Panov–Veryovkin generators are important for the analogy between the group $\pi_{1}(\mathcal{R}_{\mathcal{K}})$ and the Hopf algebra $H_{*}(\Omega\mathcal{Z}_{\mathcal{K}};\mathbf{k})$ (see Subsection 6.4 below). However, our choice of generators is more suitable for computations.
@@
For right-angled Coxeter groups $\underline{G}^{\mathcal{K}}=(\underline{\mathbb{Z}_{2}})^{\mathcal{K}}=\mathrm{RC}_{\mathcal{K}}$ we omit the subscript ${\underline{g}}$ since then $|G_{j}^{*}|=1$. Li Cai proved ([Cai21*], unpublished) that the set $\widehat{D}=\{\widehat{L}(i,J):J\subset[m],i\in\Theta(J)\}\subset\mathrm{RC}_{\mathcal{K}}^{\prime}$ generates the group $\mathrm{RC}_{\mathcal{K}}^{\prime}.$ His approach relies on a study of the Davis complex [Dav08, Chapter 7], a contractible cubical complex which is the universal covering of $\mathcal{R}_{\mathcal{K}}.$
@@ 4.2. Reduction to the distinguished elements
    in $\mathcal{K}_{J},$ i.e., a sequence of edges $\{i_{t},i_{t+1}\}\in\mathcal{K}_{J}$. Applying the relation (4.1), we replace $L_{\underline{g}}(i_{t},J)$ with a product of $L_{\underline{g}}(i_{t+1},J)$ and generators of smaller length. On the last step we replace $L_{\underline{g}}(\max(J),J)$ with $1$ (using (4.2)). Therefore, $L_{\underline{g}}(i,J)$ is replaced with a word on elements of smaller length. By the inductive hypothesis, each of this elements can be expressed through the distinguished elements.
@@ 4.4. Examples
$\cdot$ $L(4,1245)L(3,1235)^{-1}\cdot L(5,1235)L(4,1234)^{-1}\cdot L(1,1234)L(5,2345)^{-1}\Big{)}=$
@@ 5.1. Outline of the proof
Consider the polyhedral product $B:=(\operatorname{cone}(\underline{G}),\underline{G})^{\mathcal{K}}\simeq(E\underline{G},\underline{G})^{\mathcal{K}}$. In Proposition 5.11 we give a geometric interpretation of the isomorphism $\pi_{1}(B)\cong\operatorname{Cart}(\underline{G},\mathcal{K})$, identifying certain loops in $B$ with elements of $\operatorname{Cart}(\underline{G},\mathcal{K})$. Then, following Li Cai [Cai21*], we represent $B$ as the union of contractible subspaces ${\mathrm{cc}}(\mathcal{K})_{\underline{g}}$ (see (5.1)) over all ${\underline{g}}\in\prod_{i=1}^{m}G_{i}$.
For a subset $Q\subset\prod_{i=1}^{m}G_{i}$, denote $B_{Q}:=\bigcup_{{\underline{h}}\in Q}{\mathrm{cc}}(\mathcal{K})_{\underline{h}}$. Then $B_{Q\sqcup\{{\underline{g}}\}}=B_{Q}\cup{\mathrm{cc}}(\mathcal{K})_{\underline{g}},$ where ${\mathrm{cc}}(\mathcal{K})_{\underline{g}}$ is contractible. In some cases the intersection $B_{Q}\cap{\mathrm{cc}}(\mathcal{K})_{\underline{g}}$ is homotopy equivalent to $|\mathcal{K}_{J}|$ for $J=\operatorname{supp}{\underline{g}}\subset[m]$ (Proposition 5.9). Hence, by an application of the van Kampen theorem (Lemma 5.1), the group $\pi_{1}(B_{Q\sqcup\{{\underline{g}}\}})$ is obtained from $\pi_{1}(B_{Q})$ by imposing $\operatorname{rank}\Pi_{1}(\mathcal{K}_{J})$ relations and then adding $\widetilde{b}_{0}(\mathcal{K}_{J})$ generators. By induction, in Theorem 5.14 we give presentations for the fundamental group of $B_{Q}$ whenever $Q\subset\prod_{i=1}^{m}G_{i}$ is an admissible subset (see Definition 5.6). A presentation of $\pi_{1}(B)$ is then obtained by passing to the limit (Lemma 5.21).
@@ 5.2. The geometric model for the classifying space
The polyhedral product $(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$ and its subspace ${\mathrm{cc}}(\mathcal{K})_{\underline{1}}$ is shown in Figure 2 for the case $\mathcal{K}=\bullet\!\!\!-\!\!\!\bullet\bullet$ and $\underline{G}=(\mathbb{Z}_{2},\mathbb{Z}_{2},\mathbb{Z}_{2})$ (left) and $\underline{G}=(\mathbb{Z}_{2},\mathbb{Z}_{3},\mathbb{Z}_{2})$ (right).
@@
Let $x\in X$ be a basepoint. For each $\alpha\in PC,$ choose a basepoint $a_{\alpha}\in A_{\alpha}$, a path $t_{\alpha}$ in $X$ from $x$ to $a_{\alpha}$, and a set of loops $\{\lambda_{i,\alpha}:i\in I_{\alpha}\}$ in $A_{\alpha}$ that generate the group $\pi_{1}(A_{\alpha},a_{\alpha})$, see Fig. 3.
@@
Now denote $T:=\bigcup_{\alpha\in PC}\overline{t_{\alpha}}.$ We have the following pushout diagram of inclusion of connected CW-complexes with the basepoint $x$:
${A\cup T\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i}$$\scriptstyle{j}$${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C\cup T\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Y.}$
@@
    The isomorphism $\Pi_{1}(A)\overset{\cong}{\longrightarrow}\pi_{1}(A\cup T,x)$ is described as follows: if $\lambda$ is a loop in $A_{\alpha}$ with the basepoint $a_{\alpha},$ then $t_{\alpha}\cdot\lambda\cdot t_{\alpha}^{-1}$ is the corresponding loop in $A\cup T$ with the basepoint $x.$ The composite map $i_{*}:\Pi_{1}(A)\cong\pi_{1}(A\cup T,x)\to\pi_{1}(X,x)$ has the same description.
  • We show that $\pi_{1}(C\cup T,x)$ is freely generated by the set of loops $\{t_{0}\cdot s_{\alpha}\cdot t_{\alpha}^{-1}:\alpha\in PC\setminus\{0\}\}$. Indeed, consider the contractible CW-subcomplex $T^{\prime}=\bigcup_{\alpha\in PC\setminus\{0\}}\overline{s_{\alpha}}$ of $C$. Then $(C\cup T)/T^{\prime}\cong C\vee\bigvee_{\alpha\in PC\setminus\{0\}}S^{1}$, where the circles are images of the loops $\{t_{0}\cdot s_{\alpha}\cdot t_{\alpha}^{-1}:\alpha\in PC\setminus\{0\}\}$ in $C\cup T$. Since $\pi_{1}(C)=1$, the group $\pi_{1}(C\cup T,x)\cong\pi_{1}((C\cup T)/T^{\prime},x)$ is freely generated by these loops by the van Kampen theorem.
@@
where $i_{*}$ is the free product of maps
$\pi_{1}(A_{\alpha},a_{\alpha})\to\pi_{1}(X,x),\quad\lambda\mapsto t_{\alpha}\cdot\lambda\cdot t_{\alpha}^{-1},$
while $j_{*}$ is the free product of composite maps
$\pi_{1}(A_{\alpha},a_{\alpha})\to\underbrace{\pi_{1}(C,a_{0})}_{=1}\to\pi_{1}(C\cup T,x),\quad\lambda\mapsto s_{\alpha}\cdot\lambda\cdot s_{\alpha}^{-1}\mapsto t_{0}\cdot s_{\alpha}\cdot\lambda\cdot s_{\alpha}^{-1}\cdot t_{0}^{-1}.$
Since $\pi_{1}(C)=1$, the map $j_{*}$ is trivial. Finally, each group $\pi_{1}(A_{\alpha},a_{\alpha})$ is generated by the set $\{\lambda_{i,\alpha}:i\in I_{\alpha}\}$, so the pushout is isomorphic to the group
$F(t_{0}\cdot s_{\alpha}\cdot t_{\alpha}^{-1}:\alpha\in PC\setminus\{0\})\,\ast\,\pi_{1}(X)/(t_{\alpha}\cdot\lambda_{i,\alpha}\cdot t_{\alpha}^{-1}:\alpha\in PC,i\in I_{\alpha}).\qed$
@@ 5.3. Subspaces of ${\mathrm{cc}}(\mathcal{K})$
We use $\mathcal{N}$ for $\mathcal{N}(J)$ throughout the proof. By the previous lemma, we have ${\mathrm{cc}}(\mathcal{K})\cap\{x_{j}=0\}=\iota_{\mathcal{K}}(|\operatorname{st}_{\mathcal{K}^{\prime}}(\{j\})|)$, so we obtain the first statement by taking the union over $j\in J$.
@@
Consider the full subcomplex of $\mathcal{K}^{\prime}$ on the vertex set $\{(I):I\in\mathcal{K},I\cap J\neq\varnothing\}$. Since $I\cap J\neq\varnothing$ if and only if $\{j\}\subset I$ for some $j\in J$, this full subcomplex coincides with $\mathcal{N}$.
@@
Hence we can define a strict deformation retraction of $|\mathcal{N}|$ onto $\beta(|\mathcal{K}_{J}|)$ as follows: for every $I\in\mathcal{K}$ such that $I\cap J\neq\varnothing,$ we move $b_{I}$ linearly to the point $b_{I\cap J}$ along the segment $|(I\cap J\subset I)|\subset|\mathcal{N}|.$ We extend this homotopy linearly onto geometric simplices of $|\mathcal{N}|$. This homotopy is well defined: if $(I_{1}\subsetneq\dots\subsetneq I_{r})\in\mathcal{N}$, then the geometric simplex $\operatorname{conv}(b_{I_{1}},\dots,b_{I_{r}})\subset|\mathcal{N}|$ is linearly deformed onto the geometric simplex $\operatorname{conv}(b_{I_{1}\cap J},\dots,b_{I_{r}\cap J})$; the image of the homotopy lies in $\operatorname{conv}(b_{I_{1}},\dots,b_{I_{r}},b_{I_{1}\cap J},\dots,b_{I_{r}\cap J})\subset\beta(|I_{r}|)\subset|\mathcal{N}|.$